OpenStudy (anonymous):

Point B is between A and C On Line AC . Use the given information to write an equation in terms of x. Solve the equation. Then find AB and BC, and determine whether line AB and line BC are congruent.

OpenStudy (anonymous):

12. AB= x+3 13. AB= 3x-7 BC= 2x+1 BC= 3x-1 AC = 10 AC= 16